the sum of the digits f a two-digit number is 11.if the digits at ten's place is increased by 5 and the digit at unit place is decreased by 5,the digits of the number are found to be reversed.find the original number.
Answer by lwsshak3(11628)

the sum of the digits f a two-digit number is 11.if the digits at ten's place is increased by 5 and the digit at unit place is decreased by 5,the digits of the number are found to be reversed.find the original number.
***
let u=units digit
let t=tens digit
u+t=11
t=11-u
..
10(t+5)+(u-5)=10u+t
10(11-u+5)+(u-5)=10u+11-u
10(16-u)+(u-5)=10u+11-u
160-10u+u-5=9u+11
18u=144
u=8
t=11-u=3
original number=38